# Marcus Webb

### v1

@@ Abstract
Low-rank matrix, low-rank kernel, Zolotarev numbers, Grothendieck duality
@@ 1 Introduction
In Figure 1, we set $N=100$ and compare the bounds based on polynomial interpolation of $K(x,y)=\Gamma(x+y-1/2)/\Gamma(x+y)$ at mapped Chebyshev points on $[0,N]$ (following Little–Reade little1984eigenvalues ), the bounds of Beckermann–Townsend based on displacement structure implicit in positive-definite Hankel matrices (beckermann2019bounds, , Cor. 5.5), the bounds and rational interpolants presented in this work, and the singular values of the matrix (the best possible). Although the existing error bounds from the literature decay exponentially, they are extremely pessimistic compared to Theorem 1.1.
@@
While abstract, this formulation unifies continuous kernels and discrete matrices and tensors in a single framework. For continuous kernels we take $D$ and $E$ to be continua with $\lambda$ and $\mu$ their respective Lebesgue measures, whereas for a finite dimensional matrix we take
@@
densely defined for $h$ that is analytic in some open set $U\supset F$ and $\Gamma\subset U\setminus F$ is any finite union of rectifiable Jordan curves with winding number $\mathrm{Ind}_{\Gamma}(F)=1$, and the Bergman space $H_{\nu}^{p}(F)$ is the closure in $L^{p}_{\nu}(F)$ of functions that are analytic on $F$.
It is the appearance of the novel numbers $Z_{n}(L_{\mu}^{2}(E),L_{\nu}^{p}(F))$ that demonstrates that $\mathcal{K}$ is nearly-low-rank, because it must decay exponentially with $n$ at a rate that depends on how well separated $E$ and $F$ are. Indeed, in Lemma 3.2 we show that
@@
is the $n$th (classical) Zolotarev number for $E$ and $F$ zolotarev1877application ; wilber2021computing . It is known that $\lim_{n\to\infty}\left(Z_{n}(E,F)\right)^{1/n}=\exp\left(-\frac{1}{\mathrm{cap}(E,F)}\right)=:\rho(E,F)$, where $\mathrm{cap}(E,F)$ is the condenser capacity of a condenser with plates $E$ and $F$, a conformally invariant measure of the separation of subsets of $\widehat{\mathbb{C}}$ and $\rho(E,F)$ is known as the conformal modulus gonvcar1969zolotarev . In many cases $Z_{n}(E,F)=\mathcal{O}(\rho(E,F)^{-n})$ rubin2022bounding . The bound (1.7) is not used in the theorem because there are interesting cases in which $\left\|(y-z)^{-1}\right\|_{L_{\mu}^{2}(E)\to L_{\nu}^{p}(F)}=\infty$ but $Z_{n}(L_{\mu}^{2}(E),L_{\nu}^{p}(F))$ decays at the same rate as $Z_{n}(E,F)$ (see Lemma 3.5). Furthermore, in Section 6.2 we provide an example to show that the low-rank approximations derived from $Z_{n}(E,F)$ can be worse than those derived more directly from $Z_{n}(L_{\mu}^{2}(E),L_{\nu}^{p}(F))$ (by a factor of 50 in that instance).
Equation (1.6) introduces a new operator $\mathcal{K}^{\prime}$, which translates the singularities that $K(x,\cdot)$ has in $F$ to an action (such as evaluation, differentiation, or convolution) on functions defined on $F$. For an explicit construction of suitable contours $\Gamma$ for general closed sets $F$ and open neighbourhoods $U$, see (ahlfors1979complex, , Sec. 4, Ch. 4). In Section 4 we discuss how $\mathcal{K}^{\prime}$ can be interpreted in terms of Grothendieck’s theory of duality for the function space $H(F;\,L_{\lambda}^{2}(D))$ of weakly analytic functions with values in $L_{\lambda}^{2}(D)$ grothendieck1953certains . In Section 5, we show that the low-rank kernel $K_{n}$ is a rational interpolant of $K(x,\cdot)$, whose interpolation nodes and poles are the roots and poles of the optimal $\phi$ in Equation (1.5). This leads to practical algorithms for the construction of the low-rank approximant by rational interpolation, limited only by the identification of $\phi$. Section 6 provides several examples of how to apply Theorem 1.1, compares the new bounds with the optimal errors, and shows that the computable rational approximations achieve the bounds. Section 7 provides some concluding remarks, and indicates future directions and impact.
@@ 2 Proof of the Main Theorem
Please observe Figure 2 to make clear the various sets, contours and points in what follows. By Cauchy’s integral formula, for any $x\in D$, $y\in E$,
@@
where $\tilde{\Gamma}$ is any suitable contour with winding number $\mathrm{Ind}_{\tilde{\Gamma}}(E)=-1$ (i.e. it winds around once clockwise). Since $K(x,\cdot)$ is analytically continuable to $F^{\prime}$, for any open set $U$ containing $F$, the contour can be deformed to $\Gamma\subset U\setminus F$ so that $\Gamma$ has winding number $\mathrm{Ind}_{\Gamma}(F)=1$, without changing the value of the integral. If $F$ is a compact set, so that $\infty\in\widehat{\mathbb{C}}\setminus(E\cup F)$ then we must take care when deforming over the point $\infty$. The decomposition $K(x,\xi)(y-\xi)^{-1}=a(x,y)\xi^{-1}+b(x,y,\xi)\xi^{-2}$ valid for $|\xi|>R$ for sufficiently large $R$, where $b(x,y,\xi)$ is bounded as $\xi\to\infty$ can be used to show that the contribution of the contour $\Gamma_{R}=\{\xi:|\xi|=R\}$ surrounding $\infty$ is zero.
@@
Figure 2: A diagram showing the relationship between the sets, contours and points occurring in the proof of Theorem 1.1. Note that $\infty$ could also be located within $F$, unlike in this depiction.
@@
Let $h\in H(F)$. We need to show that $f=\mathcal{K}^{\prime}[h]\in C(D)$. Since $F$ is a closed subset of $\widehat{\mathbb{C}}$, there exists an open set $U$ containing $F$ such that $h\in H(U)$. Further, for any $x\in D$, $K(x,\cdot)\in C(F^{\prime})$, so both $K(x,\xi)$ and $h(\xi)$ are bounded for $\xi\in\Gamma$. Therefore, $f(x)$ is defined for all $x\in D$. Now consider a sequence $\{x_{k}\}_{k=1}^{\infty}\subset D$ converging to $x\in D$, and associated with that, consider the sequence of functions $K(x_{k},\cdot)\in C(F^{\prime})$. Since $K\in C(D\times F^{\prime})$ and by the Heine–Cantor Theorem (continuous functions on compact sets are uniformly continuous), we have that $K(x_{k},\cdot)\to K(x,\cdot)$ uniformly on compact subsets of $F^{\prime}$. Since $\Gamma$ is a compact subset of $F^{\prime}$, we have $\int_{\Gamma}K(x_{k},\xi)h(\xi)\,\mathrm{d}\xi\to\int_{\Gamma}K(x,\xi)h(\xi)\,\mathrm{d}\xi$ as $k\to\infty$. Therefore, $f$ is continuous.
@@ 3 Cauchy–Zolotarev numbers
Some elegant special cases can be derived from Equation (3.3) and Lemma 3.4. For example, with the notation of Theorem 3.3,
@@ 4.2 Grothendieck duality
where $h$ is analytic in an open set $U$ containing $F$ and $\Gamma\subset U\setminus F$ is any finite union of rectifiable Jordan curves with winding number $\mathrm{Ind}_{\Gamma}(F)=1$ silva1950funcoes . The inverse operator is $\Phi^{-1}[\ell](y)=\ell[(y-z)^{-1}]$, where $\ell\in H^{\prime}(F)$ and $y\in F^{\prime}$. If $F$ is a smooth contour then $\Phi[g]$ depends only on the jump of $g\in H(F^{\prime})$ across $F$, and $\Phi[g]$ can be interpreted in a manner more familiar to Numerical Analysts, as a hyperfunction sato1959theory .
@@ 5 Computation by rational interpolation
$K_{n}(x,y)=\frac{1}{2\pi\mathrm{i}}\int_{\Gamma}\left(1-\frac{\phi(y)}{\phi(\xi)}\right)\,\frac{K(x,\xi)}{y-\xi}\,\mathrm{d}\xi,$ (5.1)
where $\phi\in\mathcal{R}_{n,n}$ has zeros $q_{1},\ldots,q_{\ell}$ and poles $p_{1},\ldots,p_{m}$, where $\ell,m\leq n$. This formula bears a resemblance to the Hermite integral formula for rational interpolation with prescribed poles.
@@
Let $\Omega$ be a bounded open subset of $\mathbb{C}$ whose boundary $\Gamma$ is a finite union of rectifiable Jordan curves. Let $g\in H(\Omega)$, $q_{1},\ldots,q_{\ell}\in\Omega$ and $p_{1},\ldots,p_{m}\in\mathbb{C}\setminus\{q_{1},\ldots,q_{\ell}\}$. There exists a unique $r\in\mathcal{R}_{\ell-1,m}$ that interpolates $g$ at $\{q_{1},\ldots,q_{\ell}\}$ (in the Hermite sense if points are repeated) and has poles at $\{p_{1},\ldots,p_{m}\}$, which for $y\in\Omega$ satisfies
$r(y)=\frac{1}{2\pi\mathrm{i}}\int_{\Gamma}\left(1-\frac{\phi(y)}{\phi(\xi)}\right)\,\frac{g(\xi)}{\xi-y}\,\mathrm{d}\xi.$
Theorem 5.1 indeed implies, after changing the orientation of the contour and reversing the sign of the $\xi-y$ term, that the low-rank kernel given in the proof of Theorem 1.1 is a rational interpolant of $K(x,\cdot)$, with roots and poles that are independent of $x$. The roots and poles in the proof are chosen so that $\phi$ is optimal for $Z_{n}(L_{\mu}^{2}(E),L_{\nu}^{p}(F))$. If the roots of $\phi$ are distinct, then
@@
In this case, $q_{1},\ldots,q_{n}$ are distinct, so formula (5.3) can be used to compute the kernel $K_{n}$. However, for other sets $E$ and $F$, the roots of the optimal $\phi$ may not be distinct. In such cases, the rational interpolant is of Hermite type and requires partial derivatives of $K$ with respect to $y$. Indeed, for $R>1$, the optimal Zolotarev rational function for $Z_{n}(\{|z|\leq 1\},\{|z|\geq R\})$ is $\phi(\xi)=\xi^{n}$.
@@ 6.1 Cauchy matrices and tensors
where $\Gamma$ is any suitable contour within the region of analyticity of $h$, winding once around D-D. It is clear that $\mathcal{K}^{\prime}$ merely reflects $h$ across the origin, so if we take $\mathrm{d}\nu(z)=\mathrm{d}\lambda(-z)$, then $\|\mathcal{K}^{\prime}\|_{H_{\nu}^{2}(-D)\to L^{2}_{\lambda}(D)}=1$. Therefore, by Theorem 1.1,
@@ 7 Conclusion
The core of the proof is the operator decomposition $\mathcal{K}=\mathcal{K}^{\prime}\,\mathcal{C}$, where $\mathcal{C}:L_{\mu}^{2}(E)\to H(F)$ is a Cauchy transform operator and $\mathcal{K}^{\prime}:H(F)\to C(D)$ is the Grothendieck dual operator. This reveals a latent displacement structure in $\mathcal{K}$, since $\mathcal{C}$ satisfies the displacement equation,
@@
The duality of $H(F)$ and $H(F^{\prime})$ discussed in Section 4 directly impacts Potential Theory and approximation of analytic functions. The standard potential-theoretic methods of bounding the integral in the Hermite integral formula for a function $f$ with singularities within the set $F$ lead to an error bound for each contour $\Gamma$, which for fixed $n$ often tends to infinity as $\Gamma$ approaches $\partial F$ levin2006potential . Grothendieck duality takes $\Gamma$ out of the equation (after all, said equation is invariant under the choice of $\Gamma$), resulting in a single error bound involving a norm of $f$, interpreted as an element of $H^{\prime}(F)$, multiplied by a Cauchy–Zolotarev number (or its polynomial analogue). This is the topic of a future paper.
The choice of interpolation points and poles is ripe for investigation. More research on computation of Zolotarev rational functions associated with continuous and, more importantly, discrete subsets of $\widehat{\mathbb{C}}$ is needed, since Trefethen and Wilber’s algorithm is not 100% reliable trefethen2025computation . Figure 4 highlights the potential of direct computation of the optimal rational functions for Cauchy-Zolotarev numbers. When $F$ is unknown, the best approach to finding good interpolation points and poles may involve the AAA algorithm nakatsukasa2018aaa applied to $K$. Developing a practical interpolation scheme in this setting could have significant consequences, impacting any algorithm that employs low-rank approximation of analytic kernels.

### v2

@@ Abstract
Low-rank matrix, low-rank kernel, interpolative decomposition, Zolotarev numbers
@@ 1 Introduction
In Figure 1, we set $N=100$ and compare the bounds based on polynomial interpolation of $K(x,y)=\Gamma(x+y-1/2)/\Gamma(x+y)$ at mapped Chebyshev points on $[0,N]$ (following Little–Reade little1984eigenvalues ), the bounds of Beckermann–Townsend based on displacement structure implicit in positive-definite Hankel matrices (beckermann2019bounds, , Cor. 5.5) (which ignore any analyticity of the kernel), the bounds and rational interpolants presented in this work, and the singular values of the matrix (the best possible). Although the existing error bounds from the literature decay exponentially, they are extremely pessimistic compared to Theorem 1.1.
@@
While abstract, this formulation is carefully considered and unifies continuous kernels and discrete matrices and tensors in a single framework. For continuous kernels we take $D$ and $E$ to be continua with $\lambda$ and $\mu$ their respective Lebesgue measures, whereas for a finite dimensional matrix we take
@@
defined for any $h$ that is analytic on $F$, which implies that $h$ is analytic in some open set $U\supset F$, and $\Gamma\subset U\setminus F$ is any finite sum of rectifiable Jordan curves with winding numbers satisfying $\mathrm{Ind}_{\Gamma}(F)-\mathrm{Ind}_{\Gamma}(E)=1$. The Bergman space $H_{\nu}^{p}(F)$ is the closure in $L^{p}_{\nu}(F)$ of functions that are analytic on $F$.
It is the appearance of the numbers $Z_{n}(L_{\mu}^{2}(E),L_{\nu}^{p}(F))$ that demonstrates that $\mathcal{K}$ is nearly-low-rank, because these numbers must decay exponentially with $n$ at a rate that depends on how well separated $E$ and $F$ are. Indeed, in Lemma 3.2 we show that
@@
is the $n$th (classical) Zolotarev number for $E$ and $F$ zolotarev1877application ; wilber2021computing . It is known that $\lim_{n\to\infty}\left(Z_{n}(E,F)\right)^{1/n}=\exp\left(-\frac{1}{\mathrm{cap}(E,F)}\right)=:\rho(E,F)$, where $\mathrm{cap}(E,F)$ is the condenser capacity of a condenser with plates $E$ and $F$ and $\rho(E,F)$ is known as the conformal modulus, both conformally invariant measures of the separation of subsets of $\widehat{\mathbb{C}}$ gonvcar1969zolotarev . Thus, for a given $K$, you want to take $E$ and $F$ as small as possible. In many cases $Z_{n}(E,F)=\mathcal{O}(\rho(E,F)^{-n})$ rubin2022bounding . The bound (1.7) is not used in the theorem because there are interesting cases in which $\left\|(y-z)^{-1}\right\|_{L_{\mu}^{2}(E)\to L_{\nu}^{p}(F)}=\infty$ but $Z_{n}(L_{\mu}^{2}(E),L_{\nu}^{p}(F))$ decays at the same rate as $Z_{n}(E,F)$ (see Lemma 3.5). Section 6.2 contains an example showing that low-rank approximations derived from $Z_{n}(E,F)$ can be worse than those derived more directly from $Z_{n}(L_{\mu}^{2}(E),L_{\nu}^{p}(F))$.
Equation (1.6) introduces a new operator $\mathcal{K}^{\prime}$, which translates the singularities that $K(x,\cdot)$ has in $F$ to an action (such as evaluation, differentiation, or convolution) on functions defined on $F$. For an explicit construction of suitable contours $\Gamma$ for general compact sets and open neighbourhoods, see (ahlfors1979complex, , Sec. 4, Ch. 4). In Section 4 we discuss how $\mathcal{K}^{\prime}$ can be interpreted in terms of Grothendieck’s theory of duality for the function space $H(F;\,L_{\lambda}^{2}(D))$ of weakly analytic functions with values in $L_{\lambda}^{2}(D)$ grothendieck1953certains . In Section 5, we show that the low-rank kernel $K_{n}$ is a rational interpolant of $K(x,\cdot)$, whose interpolation nodes and poles are the roots and poles of the optimal $\phi$ in Equation (1.5). This leads to practical algorithms for the construction of the low-rank approximant by rational interpolation, limited only by the identification of $\phi$. Section 6 provides several examples of how to apply Theorem 1.1, compares the new bounds with the optimal errors, and shows that the computable rational approximations achieve the bounds. Section 7 provides some concluding remarks, and indicates future directions and impact.
@@ 2 Proof of the Main Theorem
By Cauchy’s integral formula, for any $x\in D$, $y\in E$,
@@
where $\tilde{\Gamma}\in F^{\prime}\setminus E$ is any finite sum of rectifiable Jordan curves, with winding numbers satisfying $\mathrm{Ind}_{\tilde{\Gamma}}(F)-\mathrm{Ind}_{\tilde{\Gamma}}(E)=1$ (note that this goes against the usual sign convention). If $\infty\in F$ then for any open set $U\supset F$, we have that $U^{\prime}$ is a bounded set containing $E$ and we can deform the contour to $\Gamma\subset U\setminus F$ and preserve the winding numbers. If $F$ is a compact set as in Figure 2, so that $\infty\in\widehat{\mathbb{C}}\setminus(E\cup F)$, then we must take care when deforming over the point $\infty$. The decomposition $K(x,\xi)(y-\xi)^{-1}=a(x,y)\xi^{-1}+b(x,y,\xi)\xi^{-2}$, valid for $|\xi|>R$ for sufficiently large $R$, where $b(x,y,\xi)$ is bounded as $\xi\to\infty$ can be used to show that the contribution of the contour $\Gamma_{R}=\{\xi:|\xi|=R\}$ surrounding $\infty$ is zero, leaving the only contributing contour to be $\Gamma$ with modified winding numbers (because of the dropped contour around infinity), but still satisfying $\mathrm{Ind}_{\Gamma}(F)-\mathrm{Ind}_{\Gamma}(E)=1$.
@@
Figure 2: A diagram showing the relationship between the sets, contours and points occurring in the proof of Theorem 1.1. Here $\mathrm{Ind}_{\tilde{\Gamma}}(E)=-1$, $\mathrm{Ind}_{\tilde{\Gamma}}(F)=0$, $\mathrm{Ind}_{\Gamma}(E)=0$, $\mathrm{Ind}_{\Gamma}(F)=1$. Note that $\infty$ could also be located within $F$, discussed in the text.
@@
Let $h\in H(F)$. We need to show that $f=\mathcal{K}^{\prime}[h]\in C(D)$. Since $F$ is a closed subset of $\widehat{\mathbb{C}}$, there exists an open set $U$ containing $F$ such that $h\in H(U)$. Further, for any $x\in D$, $K(x,\cdot)\in C(F^{\prime})$, so both $K(x,\xi)$ and $h(\xi)$ are bounded for $\xi\in\Gamma$. Therefore, $f(x)$ is defined for all $x\in D$. Now consider a sequence $\{x_{k}\}_{k=1}^{\infty}\subset D$ converging to $x\in D$, and associated with that, consider the sequence of functions $K(x_{k},\cdot)\in C(F^{\prime})$. Since $K\in C(D\times F^{\prime})$ and by the Heine–Cantor Theorem (continuous functions on compact sets are uniformly continuous), we have that $K(x_{k},\cdot)\to K(x,\cdot)$ uniformly on compact subsets of $F^{\prime}$. Since $\Gamma$ is a compact subset of $F^{\prime}$, we have $f(x_{k})=\int_{\Gamma}K(x_{k},\xi)h(\xi)\,\mathrm{d}\xi\to\int_{\Gamma}K(x,\xi)h(\xi)\,\mathrm{d}\xi=f(x)$ as $k\to\infty$. Therefore, $f$ is continuous on $D$.
@@ 3 Cauchy–Zolotarev numbers
Some elegant special cases can be derived from Equation (3.3) and Lemma 3.4. For example, with the notation of Theorem 3.3, we can show that $\|\mathcal{C}\|_{L^{2}(a,b)\to L^{2}(c,d)}\leq\log^{1/2}(\gamma)$, so
@@ 4.2 Grothendieck duality
where $h$ is analytic in an open set $U$ containing $F$ and $\Gamma\subset U\setminus F$ is any finite sum of rectifiable Jordan curves with winding numbers satisfying $\mathrm{Ind}_{\Gamma}(F)-\mathrm{Ind}_{\Gamma}(U^{\prime})=1$ silva1950funcoes . The inverse operator is $\Phi^{-1}[\varphi](y)=\varphi[(y-z)^{-1}]$, where $\varphi\in H^{\prime}(F)$ and $y\in F^{\prime}$. If $F$ is a smooth contour then $\Phi[g]$ depends only on the jump of $g\in H(F^{\prime})$ across $F$, and $\Phi[g]$ can be interpreted in a manner more familiar to Numerical Analysts, as a hyperfunction sato1959theory .
@@ 5 Computation by rational interpolation
$K_{n}(x,y)=\frac{1}{2\pi\mathrm{i}}\int_{\Gamma}K(x,\xi)\,\left(1-\frac{\phi(y)}{\phi(\xi)}\right)\frac{1}{y-\xi}\,\mathrm{d}\xi,$ (5.1)
where $\phi(\xi)=\prod_{i=1}^{\ell}(\xi-q_{i})/\prod_{k=1}^{m}(\xi-p_{k})$ for some $\ell,m\leq n$. This formula bears a resemblance to the Hermite integral formula for rational interpolation with prescribed poles.
@@
Let $\Omega$ be a bounded open subset of $\mathbb{C}$ whose boundary $\Gamma$ is a finite sum of rectifiable Jordan curves. Let $g\in H(\Omega)$, $q_{1},\ldots,q_{\ell}\in\Omega$ and $p_{1},\ldots,p_{m}\in\mathbb{C}\setminus\{q_{1},\ldots,q_{\ell}\}$. There exists a unique $r\in\mathcal{R}_{\ell-1,m}$ that interpolates $g$ at $\{q_{1},\ldots,q_{\ell}\}$ (in the Hermite sense if points are repeated) and has poles at $\{p_{1},\ldots,p_{m}\}$, which for $y\in\Omega$ satisfies
$r(y)=\frac{1}{2\pi\mathrm{i}}\int_{\Gamma}g(\xi)\,\left(1-\frac{\phi(y)}{\phi(\xi)}\right)\frac{1}{\xi-y}\,\mathrm{d}\xi.$
Theorem 5.1 indeed implies, after changing the orientation of the contour and reversing the sign of the $\xi-y$ term, that the low-rank kernel given in the proof of Theorem 1.1 is a rational interpolant of $K(x,\cdot)$, with interpolation nodes and poles that are independent of $x$. The roots and poles in the proof are chosen so that $\phi$ is optimal for $Z_{n}(L_{\mu}^{2}(E),L_{\nu}^{p}(F))$. If the roots of $\phi$ are distinct, then
@@
In this case, $q_{1},\ldots,q_{n}$ are distinct, so formula (5.3) can be used to evaluate the kernel $K_{n}$. However, for other sets $E$ and $F$, the roots of the optimal $\phi$ may not be distinct. In such cases, the rational interpolant is of Hermite type and requires partial derivatives of $K$ with respect to $y$. Indeed, for $R>1$, the optimal Zolotarev rational function for $Z_{n}(\{|z|\leq 1\},\{|z|\geq R\})$ is $\phi(\xi)=\xi^{n}$.
@@ 6.1 Cauchy matrices and tensors
where $\Gamma$ is any suitable contour within the region of analyticity of $h$, winding once around D-D (and zero times around $E)$. It is clear that $\mathcal{K}^{\prime}$ merely reflects $h$ across the origin, so if we take $\mathrm{d}\nu(z)=\mathrm{d}\lambda(-z)$, then $\|\mathcal{K}^{\prime}\|_{H_{\nu}^{2}(-D)\to L^{2}_{\lambda}(D)}=1$. Therefore, by Theorem 1.1,
@@ 7 Conclusion
The core of the proof is the operator decomposition $\mathcal{K}=\mathcal{K}^{\prime}\,\mathcal{C}$, where $\mathcal{C}:L_{\mu}^{2}(E)\to H(F)$ is a Cauchy transform operator and $\mathcal{K}^{\prime}:H(F)\to C(D)$ is the Grothendieck dual operator of $\mathcal{K}$. This reveals a latent displacement structure in $\mathcal{K}$, since $\mathcal{C}$ satisfies the displacement equation,
@@
The duality of $H(F)$ and $H(F^{\prime})$ discussed in Section 4 directly impacts Potential Theory and approximation of analytic functions. The standard potential-theoretic methods of bounding the integral in the Hermite integral formula for a function $f$ with singularities within the set $F$ lead to an error bound for each contour $\Gamma$, which for fixed $n$ often tends to infinity as $\Gamma$ approaches $\partial F$ levin2006potential . Grothendieck duality takes $\Gamma$ out of the equation (after all, the equation is invariant under the choice of $\Gamma$), resulting in a single error bound involving a norm of $f$, interpreted as an element of $H^{\prime}(F)$, multiplied by a Cauchy–Zolotarev number (or its polynomial analogue). This is the topic of a future paper.
The choice of interpolation points and poles is ripe for investigation. More research on computation of Zolotarev rational functions associated with continuous and, more importantly, discrete subsets of $\widehat{\mathbb{C}}$ is needed, since Trefethen and Wilber’s algorithm is not 100% reliable trefethen2025computation . Figure 4 highlights the potential of direct computation of the optimal rational functions for Cauchy-Zolotarev numbers. When $F$ is unknown, the best approach to finding good points and poles may involve the AAA algorithm nakatsukasa2018aaa applied to $K$. Developing a practical interpolation scheme in this setting could have significant consequences, impacting any algorithm employing low-rank approximation of analytic kernels.